OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of f(x)=x−lnx on the interval [1/8,8].

OpenStudy (anonymous):

I get f '(x)= 1-(1/x) and a critical point of x=0 When I evaluate f at the critical point and end points I get f(0) as DNE I get f(8)= 5.920558 f(1/8) = 2.20444 I thought these would be the absolute maximum and minimum, but apparently their not

OpenStudy (anonymous):

i tell you the method

OpenStudy (anonymous):

for critical point put f'(x)=0 1-1/x=0 x-1=0 x=1 is critical point thats the mistake you have done

OpenStudy (anonymous):

can u get it?

OpenStudy (anonymous):

oh! thank you!!

OpenStudy (anonymous):

welcome